# Believing without Seeing: Quality Scores for Contextualizing Vision-Language Model Explanations

### v1

@@ Abstract
When people query Vision-Language Models (VLMs) but cannot see the accompanying visual context (e.g. for blind and low-vision users), augmenting VLM predictions with natural language explanations can signal which model predictions are reliable. However, prior work has found that explanations can easily convince users that inaccurate VLM predictions are correct. To remedy undesirable overreliance on VLM predictions, we propose evaluating two complementary qualities of VLM-generated explanations via two quality scoring functions. We propose Visual Fidelity, which captures how faithful an explanation is to the visual context, and Contrastiveness, which captures how well the explanation identifies visual details that distinguish the model’s prediction from plausible alternatives. On the A-OKVQA and VizWiz tasks, these quality scoring functions are better calibrated with model correctness than existing explanation qualities. We conduct a user study in which participants have to decide whether a VLM prediction is accurate without viewing its visual context. We observe that showing our quality scores alongside VLM explanations improves participants’ accuracy at predicting VLM correctness by 11.1%, including a 15.4% reduction in the rate of falsely believing incorrect predictions. These findings highlight the utility of explanation quality scores in fostering appropriate reliance on VLM predictions.
@@ 1 Introduction
Prior work has explored the utility of model explanations to support user decision making Wang & Yin (2021); Bansal et al. (2021). However, natural language explanations can be misleading for inaccurate model predictions Joshi et al. (2023); Chaleshtori et al. (2024); Si et al. (2024a); Sieker et al. (2024). In Figure 1, we consider the question “What period of the day does this photo reflect?” where a VLM (incorrectly) answers “Noon” and generates a plausible explanation referencing cues like “a clock on the building” or “lighting and shadows.” This reasoning may appear highly convincing to a user without access to the image, even though the prediction is incorrect. Users are particularly susceptible to such overreliance on explanations in scenarios with such information asymmetry.
Figure 1: VLM explanations that sound plausible can mislead users. Contextualizing VLM explanations with quality scores may help users know when to rely on VLM outputs, but existing explanation qualities are not calibrated with VLM prediction accuracy. We propose evaluating two new qualities of VLM explanations: Visual Fidelity and Contrastiveness. These qualities are better calibrated with VLM correctness, and also help users make better decisions about when to believe VLM predictions.
@@
In experiments with three popular VLMs and two visual reasoning tasks, A-OKVQA Schwenk et al. (2022) and VizWiz Gurari et al. (2018), we find that our Visual Fidelity and Contrastiveness qualities better distinguish correct VLM predictions than do existing evaluation qualities Hase et al. (2020); Chen et al. (2023); Liu et al. (2023b), and are also better calibrated with VLM correctness (§3). We conduct several user studies in which participants evaluated VLM predictions without access to visual inputs (§4). We find that augmenting VLM explanations with our proposed qualities improves participants’ ability to distinguish between correct and incorrect VLM predictions. In particular, showing the product of Visual Fidelity and Contrastiveness as a single quality score leads to an 11.1% absolute improvement in user accuracy, along with a 15.4% absolute reduction in the rate of users falsely believing inaccurate predictions. Finally, we show that presenting explanation qualities via natural language descriptions instead of scores further improves user decision making accuracy (§4.2). Overall, our findings highlight the utility of evaluating the quality of VLM-generated explanations and communicating these qualities to users relying on VLM assistants.
@@ 2 Quality Scores for VLM Explanations
We define an explanation quality scoring function $S_{\varphi}(E;I,Q,a,\mathcal{A})$ [^1] For brevity, we note arguments for $S$ only when relevant. that maps a VLM-generated explanation $E$ to a score in $[0,1]$ reflecting a specific quality $\varphi$ of the trust cue of an explanation. Existing qualities for natural language explanations, primarily developed and evaluated in text-only settings, typically look at how well an explanation supports the model’s prediction Hase et al. (2020), its informativeness Chen et al. (2023), and its plausibility according to common sense priors Liu et al. (2023b). However, evaluating only these qualities is insufficient for VLM-generated explanations since they do not consider the relationship between the explanation and the visual context. Further, we find that these explanation qualities are poorly calibrated with prediction accuracy (§3) and therefore not useful for a user in determining whether to trust the VLM’s prediction or not.
To address these gaps, we introduce two image-anchored trust cues: Visual Fidelity and Contrastiveness. (Figure 1).
2.1 Visual Fidelity
We measure the visual fidelity of an explanation $E$ with respect to the image $I$ by decomposing the explanation into a set of facts about the image and individually verifying each fact (Algorithm 1). Conceptually, this is an oracle scoring function that would check each atomic claim against ground-truth visual Q&A; we approximate it with an LLM-as-judge, using a verifier $m_{\texttt{Verif}}$ (GPT-4o by default), and we audit this approximation with expert annotations and verifier swaps at Appendix˜F. Concretely, we prompt a question generation LLM $m_{\texttt{QGen}}$ to generate a set of verification questions $\mathcal{Q}^{V}$ that confirm visual details that have been mentioned in the VLM-generated explanation. The model has to generate questions such that answering “yes” confirms the presence of the detail in the image (see Table˜6 for the full LLM prompt). The number of questions generated by $m_{\texttt{QGen}}$ is dependent on the explanation and may vary across instances.
2.2 Measuring Contrastiveness
This quality captures whether the VLM’s explanation mentions all relevant visual details to identify the correct answer. The evidences and reasoning in the explanation may not identify the visual details that correctly distinguish the correct answer from alternatives. For instance, the explanation in Figure 1 fails to mention the time on the clock—the key visual cue for correctly answering the question. When such relevant cues are missed, the explanation may inadvertently support other alternative answers apart from the VLM’s prediction. Evaluating the Contrastiveness of an explanation could potentially alert users to when the model explanation fails to identify relevant visual cues.
@@ 2.2 Measuring Contrastiveness
A high Contrastiveness score indicates that the explanation identifies visual cues that not only support the VLM prediction, but also eliminate plausible alternatives.
@@ 2.3 Combining Visual Fidelity and Contrastiveness
$S_{\mathrm{avg}}(E)=\frac{S_{\mathrm{VF}}+S_{\mathrm{Contr.}}}{2};\quad S_{\mathrm{Prod}}(E)=S_{\mathrm{VF}}\times S_{\mathrm{Contr.}};\quad S_{\min}=\min(S_{\mathrm{VF}},S_{\mathrm{Contr.}}).$
@@ 3 Are Visual Fidelity and Contrastiveness Indicative of VLM Correctness?
We evaluate our quality scoring functions on their ability to indicate the accuracy of a VLM prediction.
@@
We evaluate our explanation qualities on two visual reasoning tasks: A-OKVQA Schwenk et al. (2022)—a multiple-choice VQA benchmark that requires reasoning over images using external knowledge and commonsense—and VizWiz Gurari et al. (2018)—an open-ended VQA task consisting of questions asked by blind and low-vision users about images they captured on their mobile phones. We sample 500 questions from the validation set of each task. Section˜G.2 contains details about the tasks’ pre-processing and evaluation. We do not evaluate Contrastiveness or combined quality scoring functions on VizWiz, which lacks a specified set of possible answers for each question. [^2] We tried turning this into a multiple-choice task by mining negatives from GPT-4o, however we found that the quality of negatives was not very good.
Table 1: Discriminability scores (higher is better) for different Quality Scoring Functions across three models and two visual reasoning tasks. We evaluate significance of Discriminability scores using an unpaired t-test; ∗∗∗ indicates significance at $p<0.001$, ∗∗ indicates $p<0.01$, indicates $p<0.05$.
A-OKVQA VizWiz
LLaVA Qwen2.5 GPT-4o LLaVA Qwen2.5 GPT-4o Average
Simulatability $0.106^{**\phantom{*}}$ $247^{***} 0.0$$98^{\phantom{***}} 0.0$$088^{*\phantom{**}}$ $0.158^{**}$ $130^{*} 0.0$940.094
Informativeness $18^{\phantom{***}} 0.000$$0^{\phantom{***}} 0.0$$69^{\phantom{***}} 0.0$$93^{*\phantom{**}} 0.0$$007^{\phantom{**}}$ $113^{\phantom{*}} 0.0$170.017
Plausibility $043^{**\phantom{*}} 0.0$$22^{\phantom{***}} 0.0$$28^{\phantom{***}} 0.0$$04^{\phantom{***}} 0.0$$20^{\phantom{**}} 0.0$$02^{\phantom{*}} 0.0$310.031
Visual Fidelity $181^{***} 0.0$$80^{***} 0.0$$085^{***}$ $^{***} 0.0420.042^{*\phantom{*}}$ $0.024^{\phantom{*}}$ 0.115
@@
[…]
@@
We experiment with three popular vision-language models: LLaVA-v1.5-7B Liu et al. (2023a), Qwen2.5-VL-7B Bai et al. (2025), and GPT-4o Hurst et al. (2024) [^3] We use the gpt-4o-2024-05-13 checkpoint.. Answers and rationales are generated in a two-step process: the model is first prompted to predict an answer to the question based on the image, and then prompted to generate a natural language explanation for its prediction. Table˜3 contains the exact prompts we use for answer and explanation generation.
@@
We compare our proposed quality measures against three established, text-only explanation qualities, each of which also returns a score in the range $[0,1]$. Simulatability Hase et al. (2020) evaluates whether an explanation offers sufficient evidence to logically justify the model’s prediction. Informativeness Chen et al. (2023) evaluates whether the explanation introduces new information to justify a prediction beyond just re-starting the question and prediction. Commonsense Plausibility Liu et al. (2023b) evaluates whether the explanation is in accordance with commonsense knowledge of everyday situations.
@@
$\text{Disc}(S_{q})=\mathbb{E}_{\begin{subarray}{c}1\leq i\leq N\\ \text{Acc}(a_{i})=1\end{subarray}}\bigl[S_{q}(E_{i})\bigr]-\mathbb{E}_{\begin{subarray}{c}1\leq i\leq N\\ \text{Acc}(a_{i})=0\end{subarray}}\bigl[S_{q}(E_{i})\bigr].$
@@
Table 2: Expected Calibration Error (ECE, lower is better) for explanation qualities.
A-OKVQA VizWiz
LLaVA Qwen2.5 GPT-4o LLaVA Qwen2.5 GPT-4o Average
Simulatability $0.288$ $0.219$ $0.302$ $0.305$ $0.292$ $0.494$ $0.316$
Informativeness $0.332$ $0.152$ $0.228$ $0.429$ $0.168$ $0.446$ $0.293$
Plausibility $0.162$ $0.270$ $0.317$ $0.110$ $0.310$ $0.325$ $0.248$
Visual Fidelity $0.207$ $0.137$ $0.099$ $0.271$ $0.136$ $0.160$ 0.168
Contrastiveness $0.176$ $0.136$ $0.209$ 0.174
Avg(VF, Contr.) $0.109$ $0.053$ $0.090$ 0.084
Prod(VF, Contr.) $0.164$ $0.154$ $0.233$ 0.183
Min(VF, Contr.) $0.147$ $0.144$ $0.227$ 0.173

### v2

@@ Abstract
When people query Vision-Language Models (VLMs) but cannot see the accompanying visual context (e.g. for blind and low-vision users), augmenting VLM predictions with natural language explanations can signal which model predictions are reliable. However, prior work has found that explanations can easily convince users that inaccurate VLM predictions are correct. To remedy undesirable overreliance on VLM predictions, we propose evaluating two complementary qualities of VLM-generated explanations via two quality scoring functions. We propose Visual Fidelity, which captures how faithful an explanation is to the visual context, and Contrastiveness, which captures how well the explanation identifies visual details that distinguish the model’s prediction from plausible alternatives. On the A-OKVQA, VizWiz, and MMMU-Pro tasks, these quality scoring functions are better calibrated with model correctness than existing explanation qualities. We conduct a user study in which participants have to decide whether a VLM prediction is accurate without viewing its visual context. We observe that showing our quality scores alongside VLM explanations improves participants’ accuracy at predicting VLM correctness by 11.1%, including a 15.4% reduction in the rate of falsely believing incorrect predictions. These findings highlight the utility of explanation quality scores in fostering appropriate reliance on VLM predictions.
Believing without Seeing: Quality Scores for Contextualizing Vision-Language Model Explanations
Keyu He1   Tejas Srinivasan2   Brihi Joshi2 Xiang Ren2Jesse Thomason2Swabha Swayamdipta2 1Carnegie Mellon University 2University of Southern California keyuhe@cmu.edu
@@ 1 Introduction
Prior work has explored the utility of model explanations to support user decision making Wang and Yin (2021); Bansal et al. (2021). However, natural language explanations can be misleading for inaccurate model predictions Joshi et al. (2023); Chaleshtori et al. (2024); Si et al. (2024a); Sieker et al. (2024). In Figure 1, we consider the question “What period of the day does this photo reflect?” where a VLM (incorrectly) answers “Noon” and generates a plausible explanation referencing cues like “a clock on the building” or “lighting and shadows.” This reasoning may appear highly convincing to a user without access to the image, even though the prediction is incorrect. Users are particularly susceptible to overreliance on explanations in scenarios with such information asymmetry.
Figure 1: VLM explanations that sound plausible can mislead users. Contextualizing VLM explanations with quality scores may help users know when to rely, but existing explanation qualities are not calibrated with VLM prediction accuracy. We propose Visual Fidelity and Contrastiveness to contextualize these explanations, which are better calibrated with correctness, and also help users make better decisions about when to believe VLM predictions.
@@
In experiments with three popular VLMs and three visual reasoning tasks, A-OKVQA Schwenk et al. (2022), VizWiz Gurari et al. (2018), and MMMU-Pro Yue et al. (2025), we find that our Visual Fidelity and Contrastiveness qualities better distinguish correct VLM predictions than do existing evaluation qualities Hase et al. (2020); Chen et al. (2023); Liu et al. (2023b), and are also better calibrated with VLM correctness (§3). We conduct several user studies in which participants evaluated VLM predictions without access to visual inputs (§4). We find that augmenting VLM explanations with our proposed qualities improves participants’ ability to distinguish between correct and incorrect VLM predictions. In particular, showing the product of Visual Fidelity and Contrastiveness as a single quality score leads to an 11.1% absolute improvement in user accuracy, along with a 15.4% absolute reduction in the rate of users falsely believing inaccurate predictions. Finally, we show that presenting explanation qualities via natural language descriptions instead of scores further improves user decision making accuracy (§4.2). Overall, our findings highlight the utility of evaluating the quality of VLM-generated explanations and communicating these qualities to users relying on VLM assistants.
@@ 2 Quality Scores for VLM Explanations
We define an explanation quality scoring function $S_{\varphi}(E;I,Q,a,\mathcal{A})$ [^1] For brevity, we note arguments for $S$ only when relevant. that maps a VLM-generated explanation $E$ to a score in $[0,1]$ reflecting a specific quality $\varphi$ of the reliability of an explanation. Existing qualities for natural language explanations, primarily developed and evaluated in text-only settings, typically look at how well an explanation supports the model’s prediction Hase et al. (2020), its informativeness Chen et al. (2023), and its plausibility according to common sense priors Liu et al. (2023b). However, evaluating only these qualities is insufficient for VLM-generated explanations since they do not consider the relationship between the explanation and the visual context. Further, we find that these explanation qualities are poorly calibrated with prediction accuracy (§3) and therefore not useful for a user deciding whether to trust the VLM’s prediction or not.
To address these gaps, we introduce two novel trust cues: Visual Fidelity and Contrastiveness. (Figure 1).
2.1 Visual Fidelity
@@
We measure the visual fidelity of an explanation $E$ with respect to the image $I$ by decomposing the explanation into a set of facts about the image and individually verifying each fact (Algorithm 1). Conceptually, this is an oracle scoring function that would check each atomic claim against ground-truth visual Q&A; we approximate it with an LLM-as-judge, using a verifier $m_{\texttt{Verif}}$ (GPT-4o by default), and we audit this approximation with expert annotations and verifier swaps at Appendix˜J. Concretely, we prompt a question generation LLM $m_{\texttt{QGen}}$ to generate a set of verification questions $\mathcal{Q}^{V}$ that confirm visual details that have been mentioned in the VLM-generated explanation. The model has to generate questions such that answering “yes” confirms the presence of the detail in the image (see Table˜7 for the full LLM prompt). The number of questions generated by $m_{\texttt{QGen}}$ depends on the explanation and may vary across instances.
2.2 Measuring Contrastiveness
Contrastiveness captures whether the VLM’s explanation mentions all relevant visual details to identify the correct answer. The evidences and reasoning in the explanation may not identify the visual details that correctly distinguish the correct answer from alternatives. For instance, the explanation in Figure 1 fails to mention the time on the clock—the key visual cue for correctly answering the question. When such relevant cues are missed, the explanation may inadvertently support other alternative answers apart from the VLM’s prediction. Evaluating the Contrastiveness of an explanation could potentially alert users to when the model explanation fails to identify relevant visual cues.
@@ 2.2 Measuring Contrastiveness
A high Contrastiveness score indicates the explanation identifies visual cues that not only support the prediction but also eliminate plausible alternatives.
@@ 2.3 Combining Visual Fidelity and Contrastiveness
$S_{\mathrm{avg}}(E)=\frac{S_{\mathrm{VF}}+S_{\mathrm{Contr.}}}{2};$
$\quad S_{\mathrm{Prod}}(E)=S_{\mathrm{VF}}\times S_{\mathrm{Contr.}};$
$\quad S_{\min}=\min(S_{\mathrm{VF}},S_{\mathrm{Contr.}}).$
@@ 3 Are Visual Fidelity and Contrastiveness Indicative of VLM Correctness?
We evaluate quality scoring functions on their ability to indicate the accuracy of a VLM prediction.
@@
We evaluate our explanation qualities on three visual reasoning tasks: A-OKVQA Schwenk et al. (2022)—a multiple-choice VQA benchmark that requires reasoning over images using external knowledge and commonsense; VizWiz Gurari et al. (2018)—an open-ended VQA task consisting of questions asked by blind and low-vision users about images they captured on their mobile phones; and MMMU-Pro Yue et al. (2025)—a massive multi-discipline multimodal understanding and reasoning benchmark. We sample 500 questions from the validation set of each task. Section˜K.2 contains details about the tasks’ pre-processing and evaluation. For VizWiz, which lacks a specified set of possible answers, we generate hard negatives using an LLM to enable Contrastiveness evaluation. These negatives are used for post-hoc evaluation and are not provided to the model during answer generation. [^2] We discuss the specific challenges and results of this open-set evaluation in Appendix F and validate the strategy via ablation in Appendix G.
Discriminability (higher is better)
A-OKVQA VizWiz MMMU-Pro Overall Avg
LLaVA Qwen2.5 GPT-4o Avg LLaVA Qwen2.5 GPT-4o Avg LLaVA Qwen2.5 GPT-4o Avg
Simulatability $0.106^{**\phantom{*}}$ $247^{***} 0.0$$098^{\phantom{***}}$ $0.150$ $0.088^{*\phantom{**}}$ $0.158^{**}$ $130^{*} 0.0$390.039 $0.059^{\phantom{**}}$ $195^{***} 0.0$$079^{**\phantom{*}}$ $0.111$ $0.100$
Informativeness $18^{\phantom{***}} 0.000$$0^{\phantom{***}} 0.0$$69^{\phantom{***}} 0.0$170.017 $093^{*\phantom{**}} 0.0$$007^{\phantom{**}}$ $113^{\phantom{*}} 0.0$$09\phantom{-} 0.0$$08^{\phantom{**}} 0.0$$06^{\phantom{***}} 0.0$$11^{\phantom{***}} 0.0$030.003 $0.004$
Plausibility $043^{**\phantom{*}} 0.0$$22^{\phantom{***}} 0.0$$28^{\phantom{***}} 0.0$310.031 $004^{\phantom{***}} 0.0$$20^{\phantom{**}} 0.0$$02^{\phantom{*}} 0.0$$06\phantom{-} 0.0$$06^{\phantom{**}} 0.0$$12^{\phantom{***}} 0.0$$04^{\phantom{***}} 0.0$070.007 $0.011$
Visual Fidelity $181^{***} 0.0$$80^{***} 0.0$$085^{***}$ $0.115$ $^{***} 0.0420.042^{*\phantom{*}}$ $0.024^{\phantom{*}}$ 0.102 $169^{**} 0.0$$38^{\phantom{***}} 0.0$$^{***} 0.0990.0$99 $0.106$
Contrastiveness $0.243^{***}$ $0.283^{***}$ $0.248^{***}$ $0.258$ $041^{\phantom{***}} 0.0$$54^{\phantom{**}} 0.0$$97^{\phantom{*}} 0.0$$01\phantom{-} 0.0$$047^{\phantom{**}}$ $54^{***} 0.101$$^{***} 0.101$0.101 $0.119$
Avg(VF, Contr.) $0.212^{***}$ $0.181^{***}$ $0.166^{***}$ $0.187$ $141^{***} 0.0$$48^{\phantom{**}} 0.0$$37^{\phantom{*}} 0.0$510.051 $108^{**} 0.0$$96^{***} 0.0$$096^{***}$ $0.100$ $0.112$
@@
Min(VF, Contr.) $0.295^{***}$ $0.298^{***}$ $0.255^{***}$ $\underline{0.283}$ $165^{***} 0.0$$77^{\phantom{**}} 0.0$$84^{\phantom{*}} 0.0$530.053 $0.082^{**}$ $0.127^{***}$ $0.124^{***}$ $\underline{0.111}$ $\underline{0.149}$
Table 1: Discriminability (Disc., higher is better) for different explanation quality scoring functions across three models and three visual reasoning tasks. Significance is evaluated using an unpaired t-test: ∗∗∗ $p<0.001$, ∗∗ $p<0.01$, $p<0.05$. Averages are not evaluated for significance. Notably, VizWiz–GPT-4o exhibits significantly lower and predominantly negative discriminability scores across all metrics; see Appendix˜I for a discussion of the underlying causes.
Expected Calibration Error (lower is better)
A-OKVQA VizWiz MMMU-Pro Overall Avg
LLaVA Qwen2.5 GPT-4o Avg LLaVA Qwen2.5 GPT-4o Avg LLaVA Qwen2.5 GPT-4o Avg
Simulatability $0.288$ $0.219$ $0.302$ $0.270$ $0.305$ $0.292$ $0.494$ $0.364$ $0.446$ $0.418$ $0.357$ $0.407$ $0.347$
Informativeness $0.332$ $0.152$ $0.228$ $0.237$ $0.429$ $0.168$ $0.446$ $0.348$ $0.216$ $0.706$ $0.538$ $0.487$ $0.357$
Plausibility $0.162$ $0.270$ $0.317$ $0.250$ $0.110$ $0.310$ $0.325$ $0.248$ $0.318$ $0.391$ $0.272$ $0.327$ $0.275$
Visual Fidelity $0.197$ $0.137$ $0.099$ $\underline{0.144}$ $0.271$ $0.139$ $0.160$ $\underline{0.190}$ $0.558$ $0.535$ $0.412$ $0.501$ $0.279$
Contrastiveness $0.176$ $0.136$ $0.210$ $0.174$ $0.294$ $0.292$ $0.424$ $0.337$ $0.221$ $0.276$ $0.256$ $0.251$ $0.254$
Avg(VF, Contr.) $0.109$ $0.054$ $0.090$ 0.084 $0.139$ $0.138$ $0.216$ 0.164 $0.360$ $0.389$ $0.279$ $0.343$ 0.197
Prod(VF, Contr.) $0.164$ $0.154$ $0.233$ $0.184$ $0.290$ $0.296$ $0.426$ $0.337$ $0.144$ $0.203$ $0.200$ 0.182 $0.234$
Min(VF, Contr.) $0.147$ $0.144$ $0.227$ $0.173$ $0.265$ $0.284$ $0.425$ $0.325$ $0.154$ $0.234$ $0.215$ $\underline{0.201}$ $\underline{0.233}$
Table 2: Expected Calibration Error (ECE, lower is better) for different explanation quality scoring functions across three models and three visual reasoning tasks.
@@
We experiment with three popular vision-language models: LLaVA-v1.5-7B Liu et al. (2023a), Qwen2.5-VL-7B Bai et al. (2025), and GPT-4o Hurst et al. (2024) [^3] We use the gpt-4o-2024-05-13 checkpoint.. We mainly use a two-step generation process: models first predict an answer, then generate an explanation for its prediction. Appendix˜C contains the exact prompts we use for answer and explanation generation.
@@
We compare our proposed qualities against three established text-only explanation baselines: Simulatability Hase et al. (2020), Informativeness Chen et al. (2023) and Commonsense Plausibility Liu et al. (2023b). See Appendix˜D for details.
@@
$\text{Disc}(S_{q})$ $=\mathbb{E}_{\begin{subarray}{c}1\leq i\leq N\\ \text{Acc}(a_{i})=1\end{subarray}}\bigl[S_{q}(E_{i})\bigr]$
$\quad-\mathbb{E}_{\begin{subarray}{c}1\leq i\leq N\\ \text{Acc}(a_{i})=0\end{subarray}}\bigl[S_{q}(E_{i})\bigr].$